# https://sarahweiii.github.io/meshlrm/ MeshLRM: Large Reconstruction Model for High-Quality Meshes

### v1

@@ Abstract
We propose MeshLRM, a novel LRM-based approach that can reconstruct a high-quality mesh from merely four input images in less than one second. Different from previous large reconstruction models (LRMs) that focus on NeRF-based reconstruction, MeshLRM incorporates differentiable mesh extraction and rendering within the LRM framework. This allows for end-to-end mesh reconstruction by fine-tuning a pre-trained NeRF LRM with mesh rendering. Moreover, we improve the LRM architecture by simplifying several complex designs in previous LRMs. MeshLRM’s NeRF initialization is sequentially trained with low- and high-resolution images; this new LRM training strategy enables significantly faster convergence and thereby leads to better quality with less compute. Our approach achieves state-of-the-art mesh reconstruction from sparse-view inputs and also allows for many downstream applications, including text-to-3D and single-image-to-3D generation.
Keywords:
Sparse-view reconstruction High-quality mesh Large Reconstruction Models![[Uncaptioned image]](extracted/2404.12385v1/figures/teaser.png)
@@ 1 Introduction
High-quality 3D mesh models are the core of 3D vision and graphics applications and 3D editing, rendering, and simulation tools are specifically optimized for them. 3D mesh assets are usually created manually by expert artists or alternatively reconstructed from multi-view 2D images. Traditionally this has been done with complex photogrammetry systems [55, 19, 49, 50], though recent neural approaches, such as NeRF [41], offer a simpler end-to-end pipeline through per-scene optimization. However, these neural methods typically produce volumetric representations [41, 42, 6, 33, 78, 63]; converting them into meshes requires additional optimization in post [59, 77, 70, 47]. Furthermore, both traditional and neural approaches require a large number of input images and long processing time (up to hours), limiting their applicability in time-sensitive design workflows.
@@
Our goal is efficient and accurate 3D asset creation via few-shot mesh reconstruction with direct feed-forward network inference and no per-scene optimization. We base our approach on recent large reconstruction models (LRMs) [23, 30, 73, 64] for 3D reconstruction and generation. Existing LRMs use triplane NeRFs [5] as the 3D representation for high rendering quality. While these NeRFs can be converted into meshes via a Marching Cube (MC) [39] post-processing, this leads to a significant drop in rendering quality and geometric accuracy.
@@
We address this with MeshLRM, a novel transformer-based large reconstruction model, designed to directly output high-fidelity 3D meshes from sparse-view inputs. Specifically, MeshLRM incorporates differentiable surface extraction and rendering into a NeRF-based LRM. We apply a recent Differentiable Marching Cube (DiffMC) [70] technique to extract an iso-surface from the triplane NeRF’s density field and render the extracted mesh using a differentiable rasterizer [29]. This enables end-to-end training of MeshLRM using mesh rendering losses, to optimize it to produce high-quality meshes that render realistic images.
@@
We train MeshLRM by initializing the model using volumetric NeRF rendering; since our meshing components do not introduce any new parameters they can start from these pre-trained weights. However, we find that training a mesh-based LRM with differentiable MC is still highly challenging. The primary issue lies in the (spatially) sparse gradients from the DiffMC operation, that only affect surface voxels and leave the vast empty space untouched. This leads to poor local minima for model optimization and manifests as floaters in the reconstructions. We address this with a novel ray opacity loss that ensures that empty space along all pixel rays maintains near-zero density, effectively stabilizing the training and guiding the model to learn accurate floater-free surface geometry. Our end-to-end training approach reconstructs high-quality meshes with rendering quality that surpasses NeRF volumetric rendering (see Tab. 5).
@@
Our DiffMC-based meshing technique is general and can potentially be applied to any NeRF-based LRMs for mesh reconstruction. In this work, we present a simple and efficient LRM architecture comprising of a large transformer model that directly processes concatenated multi-view image tokens and triplane tokens with purely self-attention layers to regress final triplane features for NeRF and mesh reconstruction. In particular, we simplify many complex design choices used in previous LRMs [30, 73, 64], including the removal of pre-trained DINO modules in image tokenization and replacing the large triplane decoder MLP with small two-layer ones. We train MeshLRM on the Objaverse dataset [14] with a novel low-resolution pre-training and high-resolution fine-tuning strategy. These design choices lead to a state-of-the-art LRM model with faster training and inference and higher reconstruction quality.
@@
A novel LRM-based framework that integrates differentiable mesh extraction and rendering for end-to-end few-shot mesh reconstruction.
A novel ray opacity loss for stabilizing DiffMC-based generalizable training.
An efficient LRM architecture and training strategies to enable fast and high-quality reconstruction.
@@ 2 Related Work
Mesh Reconstruction. Despite the existence of various 3D representations, meshes remain the most widely used format in industrial 3D engines. Reconstructing high-quality meshes from multi-view images is a long-standing challenge in computer vision and graphics. Classically, this is addressed via a complex multi-stage photogrammetry pipeline, integrating techniques such as structure from motion (SfM) [55, 1, 49], multi-view stereo (MVS) [19, 50], and mesh surface extraction [39, 27]. Recently, deep learning-based methods have also been proposed to address these problems for higher efficiency and accuracy [62, 58, 75, 24, 25, 11]. However, the photogrammety pipeline is not only costly, requiring dense input images and long processing time, but also often suffers from low-quality mesh rendering. Our approach enables sparse-view mesh reconstruction through direct feed-forward network inference in an end-to-end manner. We leverage neural volumetric representations and differentiable rendering, training our model to output high-quality meshes that can render realistic images.
@@
Neural Reconstruction. Neural rendering techniques have recently gained significant attention for their ability to produce high-quality 3D reconstructions for realistic rendering [80, 37, 41, 28]. Most recent methods are based on NeRF [41] and reconstruct scenes as various volumetric radiance fields [41, 6, 18, 42, 72] with per-scene rendering optimization. Though radiance fields can be converted into meshes with Marching Cubes [39], the quality of the resulting mesh is not guaranteed. Previous methods have aimed to transform radiance fields into implicit surface representations (SDFs) [63, 76, 43], enhancing mesh quality from Marching Cubes. In addition, some methods [77, 70, 59, 47] attempt to directly extract meshes from radiance fields using differentiable surface rendering. However, these methods all require dense input images and long per-scene optimization time, which our approach avoids.
@@
On the other hand, generalizable neural reconstruction has been explored, often leveraging MVS-based cost volumes [7, 38, 26] or directly aggregating multi-view features along pixels’ projective epipolar lines [79, 65, 57]. While enabling fast and few-shot reconstruction, these methods can only handle small baselines, still requiring dense images with local reconstruction in overlapping windows to model a complete object. More recently, transformer-based large reconstruction models (LRMs) [23, 30, 64, 73] have been proposed and achieved 3D NeRF reconstruction from highly sparse views. In this work, we propose a novel efficient LRM, incorporating DiffMC techniques [70], to enable high-quality sparse-view mesh reconstruction via direct feed-forward inference.
@@
3D Generation. Generative models have seen rapid progress with GANs [21] and, more recently, Diffusion Models [56] for image and video generation. In the context of 3D generation, many approaches utilize 3D or multi-view data to train 3D generative models with GANs [71, 22, 20, 5] or diffusion models [2, 73, 67, 8, 36], which are promising but limited by the scale and diversity of 3D data. DreamFusion [45] proposed to leverage the gradient of a pre-trained 2D diffusion model to optimize a NeRF for text-to-3D generation with a score distillation sampling (SDS) loss, achieving diverse generation results beyond common 3D data distributions. This approach inspired many follow-up approaches, to make the optimization faster or improve the results quality [31, 68, 60, 10, 40]. However, these SDS-based methods rely on NeRF-like per-scene optimization, which is highly time-consuming, often taking hours. More recently, a few attempts have been made to achieve fast 3D generation by utilizing pre-trained 2D diffusion models to generate multi-view images and then perform 3D reconstruction [30, 34, 35]. In particular, Instant3D [30] leverages an LRM to reconstruct a triplane NeRF from four sparse generated images. In this work, we focus on the task of sparse-view reconstruction and propose to improve and re-target a NeRF LRM to directly generate meshes with higher quality. Our model can be naturally coupled with a multi-view generator, such as the ones in [30, 53], to enable high-quality text-to-3D and single-image-to-3D generation (see Fig. 5 and Fig. 6).
@@ 3 Method
In this section, we present our MeshLRM that reconstructs high-quality meshes within 1 second. We start with describing our backbone transformer architecture (Sec. 3.1), which simplifies and improves previous LRMs. We then introduce our two-stage framework for training the model: we first (Sec. 3.2) train the model to predict NeRF from sparse input images by supervising volume renderings at novel views, followed by refining mesh surface extractions (Sec. 3.3) by performing differentiable marching cubes on the predicted density field and minimizing a surface rendering loss with differentiable rasterization.
@@ 3.1 Model Architecture
As shown in Fig. 2, we propose a simple transformer-based architecture for MeshLRM, mainly consisting of a sequence of self-attention-based transformer blocks over concatenated image tokens and triplane tokens, similar to PF-LRM [64]. In contrast to PF-LRM and other LRMs [23, 30, 73], we simplify the designs for both image tokenization and triplane NeRF decoding, leading to fast training and inference. We next describe the model in detail.
@@
Input posed image tokenization. MeshLRM adopts a simple tokenizer for posed images, inspired by ViT[16]. We convert the camera parameters for each image into Plücker ray coordinates [44] and concatenate them with the RGB pixels (3-channel) to form a 9-channel feature map. Then we split the feature map into non-overlapping patches, and linearly transform them as transformer’s input. With this process, the model does not need additional positional embedding as in ViT, since Plücker coordinates contain spatial information.
@@
Note that our image tokenizer is much simpler than previous LRMs that use a pre-trained DINO ViT [4] for image encoding. We empirically find that the pre-trained DINO ViT is unnecessary, possibly because DINO is mainly trained for intra-view semantic reasoning, while 3D reconstruction mainly requires inter-view low-level correspondences. By dropping this per-view DINO encoding, the model also enables a better flow between raw pixels and 3D-related information.
@@
Transformer. We concatenate multi-view image tokens and learnable triplane (positional) embeddings, and feed them into a sequence of transformer blocks [61], where each block is comprised of self-attention and MLP layers. We add layer normalization [3] before both layers (i.e., Pre-LN architecture), and use residual connections. This deep transformer network enables comprehensive information exchange among all the tokens, and effectively models intra-view, inter-view, and cross-modal relationships. The output triplane tokens, contextualized by all input views, are then decoded into the renderable triplane NeRF. The output image tokens are dropped. Lastly, each triplane token is unprojected with a linear layer and further unpatchified to triplane features via reshaping. 111This operator is logically the same to a shared 2D-deconvolution with stride and kernel on each plane. All the predicted triplane features are then assembled into the final triplane NeRF.
@@
Tiny density and color MLPs. Previous LRMs [23, 30, 73, 64] all use a heavy shared MLP (e.g., 9 hidden layers with a hidden dimension of 64) to decode densities and colors from triplane features. This design leads to slow rendering in training. This is a bottleneck for MeshLRM since we need to compute a dense density grid, using the MLP and triplane features, for extracting the mesh with DiffMC. Therefore, we opt to use tiny MLPs with a narrower hidden dimension of and fewer layers. In particular, we use an MLP with one hidden layer for density decoding and another MLP with two hidden layers for color decoding. Compared to a large MLP, our tiny MLPs can lead to speed-up in training without compromising quality (see Tab. 4). We use separate MLPs since the density MLP and color MLP are used separately in the Marching Cubes and surface rendering processes, respectively. We empirically find this MLP separation is necessary for DiffMC fine-tuning (described later in Sec. 3.3) as it largely improves the optimization stability which is critical to large-scale training.
@@
While being largely simplified, our transformer model can effectively transform input posed images into a triplane NeRF for density and color decoding. The density and color are then used to achieve both radiance field rendering for 1st-stage volume initialization (Sec. 3.2) and surface extraction+rendering for 2nd-stage mesh reconstruction (Sec. 3.3).

### v2

# MeshLRM: Large Reconstruction Model for High-Quality Meshes

@@ Abstract
We propose MeshLRM, a novel LRM-based approach that can reconstruct a high-quality mesh from merely four input images in less than one second. Different from previous large reconstruction models (LRMs) that focus on NeRF-based reconstruction, MeshLRM incorporates differentiable mesh extraction and rendering within the LRM framework. This allows for end-to-end mesh reconstruction by fine-tuning a pre-trained NeRF LRM using mesh rendering. Moreover, we improve the LRM architecture by simplifying several complex designs in previous LRMs. MeshLRM’s NeRF initialization is sequentially trained with low- and high-resolution images; this new LRM training strategy enables significantly faster convergence and thereby leads to better quality with less compute. Our approach achieves state-of-the-art mesh reconstruction from sparse-view inputs and also allows for many downstream applications, including text-to-3D and single-image-to-3D generation. https://meshlrm.github.io/
@@ 1 Introduction
High-quality 3D mesh models are the core of 3D vision and graphics applications that are specifically optimized for them, such as 3D editing, rendering, or simulation tools. 3D mesh assets are usually created manually by expert artists or reconstructed from multi-view 2D images. Traditionally this has been done with complex photogrammetry systems (Snavely et al., 2006; Furukawa & Ponce, 2009; Schonberger & Frahm, 2016; Schönberger et al., 2016), though recent neural representations, such as NeRF (Mildenhall et al., 2020), offer a simpler end-to-end pipeline through per-scene optimization. These neural representations are however typically volumetric (Mildenhall et al., 2020; Müller et al., 2022; Chen et al., 2022; Liu et al., 2020; Yariv et al., 2020; Wang et al., 2021a) and converting them into meshes requires additional post-optimization (Tang et al., 2022; Yariv et al., 2023; Wei et al., 2023; Rakotosaona et al., 2023). Furthermore, both traditional and neural reconstruction approaches require a large number of input images and often long processing time (up to hours), limiting their interactivity.
@@
Our goal is efficient and accurate 3D asset creation via sparse-view mesh reconstruction with direct feed-forward network inference and no post-optimization. We base our approach on recent large reconstruction models (LRMs) (Hong et al., 2024; Li et al., 2024; Xu et al., 2023; Wang et al., 2023a) for 3D reconstruction and generation. Existing LRMs use triplane NeRFs (Chan et al., 2022) as the 3D representation for high rendering quality. While these NeRFs can be converted into meshes via a Marching Cube (MC) (Lorensen & Cline, 1998) post-processing step. If done naively, this typically leads to a significant drop in rendering quality and geometric accuracy.
@@
We propose to address this with MeshLRM, a novel transformer-based large reconstruction model, designed to directly output high-fidelity 3D meshes from sparse-view inputs. Specifically, we incorporate differentiable surface extraction and rendering into a NeRF-based LRM. We apply a recent Differentiable Marching Cube (DiffMC) (Wei et al., 2023) technique to extract an iso-surface from the triplane NeRF’s density field and render the extracted mesh using a differentiable rasterizer (Laine et al., 2020). This lets us train MeshLRM end-to-end using a mesh rendering loss, optimizing it to produce density fields which are more compatible with the Marching Cube step, leading to more realistic and high-quality meshes.
@@
We train MeshLRM by initializing the model using volumetric NeRF rendering; since our meshing components do not introduce any new parameters these pre-trained weights are a good starting point for our fine-tuning. We however find that fine-tuning a mesh-based LRM with differentiable MC (DiffMC) remains challenging. The primary challenge lies in the (spatially) sparse gradients from the DiffMC operation, that only affect surface voxels and leave the vast empty space untouched. This leads to poor local minima for model optimization and manifests as floaters in the reconstructions. We address this with a novel ray opacity loss that ensures that empty space along all pixel rays maintains near-zero density, effectively stabilizing the training and guiding the model to learn accurate floater-free surface geometry. Our end-to-end training approach reconstructs high-quality meshes with rendering quality that surpasses NeRF volumetric rendering (as shown in Tab. 5).
@@
Our approach leverages a simple and efficient LRM architecture consisting of a large transformer model that directly processes concatenated multi-view image tokens and triplane tokens with self-attention layers to regress final triplane features that we use for NeRF and mesh reconstruction. In particular, we simplify many complex design choices used in previous LRMs (Li et al., 2024; Xu et al., 2023; Wang et al., 2023a), including the removal of pre-trained DINO modules in image tokenization and replacing the large triplane decoder MLP with small two-layer ones. We train MeshLRM on the Objaverse dataset (Deitke et al., 2023) with a progressive resolution strategy with low-resolution early training and high-resolution fine-tuning. These design choices lead to a state-of-the-art LRM model with faster training and inference and higher reconstruction quality.
@@
A LRM-based framework that integrates differentiable mesh extraction and rendering for end-to-end few-shot mesh reconstruction.
A ray opacity loss for improved DiffMC-based training stability.
A simplified LRM architecture and enhanced training strategies for fast and high-quality reconstruction.
We benchmark MeshLRM for 3D reconstruction (on synthetic and real datasets) and 3D generation (in combination with other multi-view generation methods). Fig. 1 showcases high-quality mesh outputs from MeshLRM that were each reconstructed in under one second.
@@ 2 Related Work
Mesh Reconstruction. Despite the existence of various 3D representations, meshes remain the most widely used format in industrial 3D engines. Reconstructing high-quality meshes from multi-view images is a long-standing challenge in computer vision and graphics. Classically, this is addressed via a complex multi-stage photogrammetry pipeline, integrating techniques such as structure from motion (SfM) (Snavely et al., 2006; Agarwal et al., 2011; Schonberger & Frahm, 2016), multi-view stereo (MVS) (Furukawa & Ponce, 2009; Schönberger et al., 2016), and mesh surface extraction (Lorensen & Cline, 1998; Kazhdan et al., 2006). Recently, deep learning-based methods have also been proposed to address these problems for higher efficiency and accuracy (Vijayanarasimhan et al., 2017; Tang & Tan, 2018; Yao et al., 2018; Huang et al., 2018; Im et al., 2019; Cheng et al., 2020). However, the photogrammetry pipeline requires dense input images and long processing time, aiming at exact object reconstruction, but often suffers from challenging calibration or appearance effects on the captures object. As opposed to this classic many-image acquisition process, our approach enables sparse-view mesh reconstruction through direct feed-forward network inference in an end-to-end manner.
@@
Neural Reconstruction. Neural rendering techniques have recently gained significant attention for their ability to produce high-quality 3D reconstructions for realistic rendering (Zhou et al., 2018; Lombardi et al., 2019; Mildenhall et al., 2020; Kerbl et al., 2023). Most recent methods are based on NeRF (Mildenhall et al., 2020) and reconstruct scenes as various volumetric radiance fields (Mildenhall et al., 2020; Chen et al., 2022; Fridovich-Keil et al., 2022; Müller et al., 2022; Xu et al., 2022) with per-scene rendering optimization. Though radiance fields can be converted into meshes with Marching Cubes (Lorensen & Cline, 1998), the quality of the resulting mesh is not guaranteed. Previous methods have aimed to transform radiance fields into implicit surface representations (SDFs) (Wang et al., 2021a; Yariv et al., 2021; Oechsle et al., 2021), enhancing mesh quality from Marching Cubes. In addition, some methods (Yariv et al., 2023; Wei et al., 2023; Tang et al., 2022; Rakotosaona et al., 2023) attempt to directly extract meshes from radiance fields using differentiable surface rendering. However, these methods all require dense input images and long per-scene optimization time, which our approach avoids.
@@
On the other hand, generalizable neural reconstruction has been explored, often leveraging MVS-based cost volumes (Chen et al., 2021; Long et al., 2022; Johari et al., 2022) or directly aggregating multi-view features along pixels’ projective epipolar lines (Yu et al., 2021; Wang et al., 2021b; Suhail et al., 2022). While enabling fast and few-shot reconstruction, these methods can only handle small baselines, still requiring dense images with local reconstruction in overlapping windows to model a complete object. More recently, transformer-based large reconstruction models (LRMs) (Hong et al., 2024; Li et al., 2024; Wang et al., 2023a; Xu et al., 2023) have been proposed and achieved 3D NeRF reconstruction from highly sparse views. Inspired by this line of work, we propose a more efficient LRM architecture, incorporating DiffMC techniques (Wei et al., 2023) to enable high-quality sparse-view mesh reconstruction via direct feed-forward inference.
@@
3D Generation. Generative models have seen rapid progress with GANs (Goodfellow et al., 2014) and, more recently, Diffusion Models (Sohl-Dickstein et al., 2015) for image and video generation. In the context of 3D generation, many approaches utilize 3D or multi-view data to train 3D generative models with GANs (Wu et al., 2016; Henzler et al., 2019; Gao et al., 2022; Chan et al., 2022) or diffusion models (Anciukevičius et al., 2023; Xu et al., 2023; Wang et al., 2023c; Chen et al., 2023a; Liu et al., 2023b), which are promising but limited by the scale and diversity of 3D data. DreamFusion (Poole et al., 2022) proposed to leverage the gradient of a pre-trained 2D diffusion model to optimize a NeRF for text-to-3D generation with a score distillation sampling (SDS) loss, achieving diverse generation results beyond common 3D data distributions. This approach inspired many follow-up approaches, to make the optimization faster or improve the results quality (Lin et al., 2023; Wang et al., 2023d; Tang et al., 2023; Chen et al., 2023b; Metzer et al., 2022). However, these SDS-based methods rely on NeRF-like per-scene optimization, which is highly time-consuming, often taking hours. More recently, a few attempts have been made to achieve fast 3D generation by utilizing pre-trained 2D diffusion models to generate multi-view images and then perform 3D reconstruction (Li et al., 2024; Liu et al., 2023a; 2024b; Tochilkin et al., 2024; Tang et al., 2024; Liu et al., 2024c). In particular, Instant3D (Li et al., 2024) leverages an LRM to reconstruct a triplane NeRF from four sparse generated images. In this work, we focus on the task of sparse-view reconstruction and propose to improve and re-target a NeRF LRM to directly generate meshes with higher quality. Unlike MeshGPT (Siddiqui et al., 2024), which sequentially predicts triangles using a transformer, we incorporate differentiable marching cubes and rasterization into the LRM framework to generate meshes. Concurrently, InstantMesh (Xu et al., 2024) makes a similar attempt by combining OpenLRM He & Wang (2023) with Flexicubes Shen et al. (2023) for direct mesh generation. In contrast, we propose a novel simplified LRM architecture with a carefully designed training strategy and supervision losses, achieving superior quality over InstantMesh. Our model can be naturally coupled with a multi-view generator, such as the ones in (Li et al., 2024; Shi et al., 2023a), to enable high-quality text-to-3D and single-image-to-3D generation (see Fig. 7 and Fig. 6).
@@ 3 Method
We present MeshLRM enabling the reconstruction of high-quality meshes in under 1 second. We start with describing our backbone transformer architecture (Sec. 3.1), which simplifies and improves previous LRMs. We then introduce our two-stage framework for training the model: we first (Sec. 3.2) train the model to predict NeRF from sparse input images by supervising volume renderings at novel views, followed by refining the model for mesh surface extractions (Sec. 3.3) by performing differentiable marching cubes on the predicted density field and minimizing a surface rendering loss with differentiable rasterization.
@@ 3.1 Model Architecture
As shown in Fig. 2, we propose a simple transformer-based architecture for MeshLRM, mainly consisting of a sequence of self-attention-based transformer blocks over concatenated image tokens and triplane tokens, similar to PF-LRM (Wang et al., 2023a). In contrast to PF-LRM and other LRMs (Hong et al., 2024; Li et al., 2024; Xu et al., 2023), we simplify the designs for both image tokenization and triplane NeRF decoding, leading to fast training and inference.
@@
Input posed image tokenization. MeshLRM adopts a simple tokenizer for posed images, inspired by ViT (Dosovitskiy et al., 2020). We convert the camera parameters for each image into Plücker ray coordinates (Plücker, 1865) and concatenate them with the RGB pixels (3-channel) to form a 9-channel feature map. We then split the feature map into non-overlapping patches, and linearly transform them to input our transformer. With this process, the model does not need additional positional embedding as in ViT, since Plücker coordinates contain spatial information.
@@
Note that our image tokenizer is much simpler than previous LRMs that use a pre-trained DINO ViT (Caron et al., 2021) for image encoding. We find that the pre-trained DINO ViT is unnecessary, possibly because it is mainly trained for intra-view semantic reasoning, while 3D reconstruction requires inter-view low-level correspondences. Vision transformer models (like DINO) often lack the zero-shot ability to handle untrained higher resolutions and require extensive fine-tuning to generalize (see Fig. 8). By dropping this per-view DINO encoding, the model becomes shallower, enabling a better connection between raw pixel information and 3D-related processing.
@@
Transformer. We concatenate multi-view image tokens and learnable triplane (positional) embeddings, and feed them into a sequence of transformer blocks (Vaswani et al., 2017), where each block is comprised of self-attention and MLP layers. We add layer normalization (Ba et al., 2016) before both layers (i.e., Pre-LN architecture), and use residual connections. This deep transformer network enables comprehensive information exchange among all the tokens, and effectively models intra-view, inter-view, and cross-modal relationships. The output triplane tokens, contextualized by all input views, are then decoded into the renderable triplane NeRF while the output image tokens are dropped. More specifically, each triplane token is unprojected with a linear layer and further unpatchified to triplane features via reshaping. 111This operator is logically the same to a shared 2D-deconvolution with stride and kernel on each plane. All predicted triplane features are then assembled into the final triplane NeRF.
@@
Tiny density and color MLPs. Previous LRMs (Hong et al., 2024; Li et al., 2024; Xu et al., 2023; Wang et al., 2023a) use a heavy shared MLP (e.g., 9 hidden layers with a hidden dimension of 64) to decode densities and colors from triplane features. This design leads to slow rendering in training. This is a bottleneck for MeshLRM since we need to compute a dense density grid, using the MLP and triplane features, for extracting the mesh with DiffMC. Therefore, we opt to use tiny MLPs with a narrower hidden dimension of and fewer layers. In particular, we use an MLP with one hidden layer for density decoding and another MLP with two hidden layers for color decoding. Compared to a large MLP, our tiny MLPs lead to speed-up in training without compromising quality (see Tab. 4). We use separate MLPs since the density MLP and color MLP are used separately in the Marching Cubes and surface rendering processes, respectively. We empirically find this MLP separation largely improves the optimization stability for DiffMC fine-tuning (described in Sec. 3.3), which is critical to large-scale training.
@@
While being largely simplified, our transformer model can effectively transform input posed images into a triplane NeRF for density and color decoding. The density and color are then used to achieve both radiance field rendering for 1st-stage volume initialization (Sec. 3.2) and surface extraction & rendering for 2nd-stage mesh reconstruction (Sec. 3.3).